# SANA-Sprint: One-Step Diffusion with Continuous-Time Consistency Distillation

### v3

@@ 6 Conclusion
In this paper, we introduced SANA-Sprint, an efficient diffusion model for ultra-fast one-step text-to-image generation while preserving multi-step sampling flexibility. By employing a hybrid distillation strategy combining continuous-time consistency distillation (sCM) and latent adversarial distillation (LADD), SANA-Sprint achieves SoTA performance with 7.04 FID and 0.72 GenEval in one step, eliminating step-specific training. This unified step-adaptive model enables high-quality 1024$\times$1024 image generation in only 0.1s on H100, setting a new SoTA in speed-quality tradeoffs.
@@ Appendix A Pseudo Code for Training-Free Transformation from Flow to Trigflow
In this section, we provide a concise implementation of the transformation from a trained flow matching model to a TrigFlow model without requiring additional training. This transformation is based on the theoretical equivalence established in Appendix D. The core idea is to first convert the TrigFlow timestep $t_{\texttt{Trig}}$ to its corresponding flow matching timestep $t_{\texttt{FM}}$. Then, the input feature $\boldsymbol{x}_{\texttt{Trig}}$ is scaled accordingly to obtain $\boldsymbol{x}_{\texttt{FM}}$. The output of the flow matching model is then transformed using a linear combination to produce the final TrigFlow output. The following pseudo code implements this transformation efficiently, ensuring consistency between the two formulations.
@@ Appendix B Transformation Algorithm
We present an algorithm for training-free transformation from a flow matching model to its TrigFlow counterpart. Given a noisy sample, its corresponding TrigFlow timestep, and a pre-trained flow matching model, the algorithm computes the equivalent flow matching timestep, rescales the input, and applies a deterministic transformation to obtain the TrigFlow output. The detailed procedure is outlined in Algorithm 1.
@@ Appendix C Training Algorithm of SANA-Sprint
11:         $\boldsymbol{\psi}\leftarrow\boldsymbol{\psi}-\eta\nabla_{\boldsymbol{\psi}}\mathcal{L}_{\text{adv}}^{D}(\boldsymbol{\psi})$triangleright Discriminator step
@@
17:         $r\leftarrow\min(1,\text{Iters}/H)$triangleright Tangent warmup
18:         $\boldsymbol{g}\leftarrow-\cos^{2}(t)(\sigma_{d}\boldsymbol{F_{\theta^{-}}}-\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t})-r\cdot\cos(t)\sin(t)(\boldsymbol{x}_{t}+\sigma_{d}\frac{\mathrm{d}\boldsymbol{F_{\theta^{-}}}}{\mathrm{d}t})$triangleright JVP rearrangement
19:         $\boldsymbol{g}\leftarrow\boldsymbol{g}/(\|\boldsymbol{g}\|+c)$triangleright Tangent normalization
20:         $\mathcal{L}(\boldsymbol{\theta,\phi})\leftarrow\frac{e^{w_{\boldsymbol{\phi}}(t)}}{D}\|\boldsymbol{F_{\theta}}(\frac{\boldsymbol{x}_{t}}{\sigma_{d}},t)-\boldsymbol{F_{\theta^{-}}}(\frac{\boldsymbol{x}_{t}}{\sigma_{d}},t)-\boldsymbol{g}\|_{2}^{2}-w_{\boldsymbol{\phi}}(t)$triangleright sCM loss
@@
24:         $\mathcal{L}(\boldsymbol{\theta,\phi})\leftarrow\mathcal{L}(\boldsymbol{\theta,\phi})-\mathbb{E}_{\boldsymbol{x}_{0},s,t}\Big{[}\sum_{k}D_{\boldsymbol{\psi},k}(\boldsymbol{F}_{\boldsymbol{\theta}_{\texttt{pre}},k}(\hat{\boldsymbol{x}}_{s}^{\boldsymbol{f_{\theta}}},s,\boldsymbol{y}))\Big{]}$triangleright GAN loss
25:         $(\boldsymbol{\theta,\phi})\leftarrow(\boldsymbol{\theta,\phi})-\eta\nabla_{\theta,\phi}\mathcal{L}(\boldsymbol{\theta,\phi})$triangleright Generator step
@@
Since $(1-t_{\texttt{FM}})^{2}+t_{\texttt{FM}}^{2}$ is generally not equal to $\cos^{2}(t_{\texttt{Trig}})+\sin^{2}(t_{\texttt{Trig}})=1$ (except when $t_{\texttt{FM}}=0$ or $1$), a scale adjustment is needed. To align their scales, we introduce a scale factor function $\lambda(t_{\texttt{FM}})$ that satisfies
@@ Appendix E Full Related Work
Text-to-image generation has experienced transformative advancements in both efficiency and model design. The field gained early traction with Stable Diffusion [59], which set the stage for scalable high-resolution synthesis. A pivotal shift occurred with Diffusion Transformers (DiT)[60], which replaced conventional U-Net architectures with transformer-based designs, unlocking improved scalability and computational efficiency. Building on this innovation, PixArt-$\alpha$[61] demonstrated competitive image quality while slashing training costs to only 10.8% of those required by Stable Diffusion v1.5 [59]. Recent breakthroughs have further pushed the boundaries of compositional generation. Large-scale models like FLUX [21] and Stable Diffusion 3 [38] have scaled up to ultra-high-resolution synthesis and introduced multi-modal capabilities through frameworks such as the Multi-modal Diffusion Transformer (MM-DiT)[22]. Playground v3[39] achieved state-of-the-art image quality by seamlessly integrating diffusion models with Large Language Models (LLMs)[62], while PixArt-$\Sigma$[37] showcased direct 4K image generation using a compact 0.6B parameter model, emphasizing computational efficiency alongside high-quality outputs. Efficiency-driven innovations have also gained momentum. SANA [32] introduced high-resolution synthesis capabilities through deep compression autoencoding [63] and linear attention mechanisms, enabling deployment on consumer-grade hardware like laptop GPUs. Additionally, advancements in linear attention mechanisms for class-conditional generation [64, 65], diffusion models without attention [66, 67], and cascade structures [68, 69, 70] have further optimized computational requirements while maintaining performance. These developments collectively underscore the field’s rapid evolution toward more accessible, efficient, and versatile text-to-image generation technologies.
@@ F.1 Experimental Setup
We conduct distributed training using PyTorch’s Distributed Data Parallel (DDP) across 32 NVIDIA A100 GPUs on 4 DGX nodes. Our two-phase strategy involves fine-tuning the teacher model with dense time embedding and QK normalization at a learning rate of 2e-5 for 5,000 iterations (global batch size of 1,024), as discussed in Sec. 3.1. Then, we perform timestep distillation through the proposed framework at a learning rate of 2e-6 with a global batch size of 512 for 20,000 iterations. As Flash Attention JVP kernel support is not available in PyTorch [17], we retain Linear Attention [32] to auto-compute the JVP.
@@ F.2 More Ablations
Fig. 6 illustrates the process of timestep optimization for inference across 1, 2, and 4 steps, comparing the performance of 0.6B and 1.6B models in terms of FID (top row) and CLIP-Scores (bottom row). The optimization follows a sequential search strategy: first, we determine the optimal $t_{\text{max}}$ for 1-step inference using $\arctan(n/0.5)$, inspired by EDM [28], where $n$ is searched for the maximum timestep. Using this $t_{\text{max}}$, we then search for the intermediate timestep $t_{2nd}$ in 2-step inference. For 4-step inference, the timesteps for the first two steps are fixed to their previously optimized values, while the third ($t_{3rd}$) and fourth timesteps ($t_{4th}$) are searched sequentially. In each case, the x-axis represents the timestep being optimized at the current step, ensuring that earlier steps use their best-found values to maximize overall performance. This hierarchical approach enables efficient timestep selection for multi-step inference settings.
@@
Timestep T $[\pi/2,0.0]$ $[\arctan(200/0.5),1.3,0.0]$ $[\arctan(200/0.5),1.3,1.1,0.6,0.0]$
@@
In the sCM-only experiments, we investigate the impact of different noise distribution parameter settings on model performance. The noise distribution is defined as $t=\arctan{\left(\frac{e^{\tau}}{\sigma_{d}}\right)}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. Starting from the initial parameters $(-0.8,1.6)$ proposed in sCM [17], we experiment with various mean and standard deviation configurations to evaluate their effects. By tracking FID and CLIP-Score trends over 40k training iterations, we identify $(P_{\text{mean}},P_{\text{std}})=(0.0,1.6)$, represented by the green curve in Fig. 7, as the optimal setting. This configuration consistently reduces FID while improving CLIP-Score, resulting in superior generation quality and text-image alignment. We also observe that extreme mean values, such as $P_{\text{mean}}=0.6$ or $P_{\text{mean}}=-0.8$, lead to significant training instability and even failure in some cases. Consequently, we adopt $(0.0,1.6)$ as the default parameter setting.
@@
Generative features change with the noise level, offering structured feedback at high noise and texture-related feedback at low noise [5]. We compare the results for different mean and standard deviation settings in $t=\arctan{\left(\frac{e^{\tau}}{\sigma_{d}}\right)}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ for LADD discriminator. Building on the optimal mean and standard deviation settings $(0.0,1.6)$ identified for sCM, we further explore the best noise configuration for the LADD’s discriminator. In Fig. 8 and Tab. 8, we visualize the probability distributions of $t$ sampled under different mean and standard deviation settings, as well as the corresponding FID and CLIP-Score results when applied in the LADD loss. Based on these analyses, we identify $(-0.6,1.0)$ as the optimal setting, which achieves a more balanced feature distribution across high and low noise levels while maintaining stable training dynamics. Consequently, we adopt $(-0.6,1.0)$ as the default configuration for LADD loss.
@@ F.3 More Qualitative Results
In Fig. 9, we demonstrate the visualization capabilities of our SANA-Sprint-ControlNet, which efficiently achieves impressive results in only 0.4 seconds using a 2-step generation process, producing high-quality images at a resolution of 1024 $\times$ 1024 pixels. The visualization process begins with an input image, which is processed using a HED detection model to extract the scribe graph. This scribe graph, combined with a given prompt, is used to generate the corresponding image in the second column. The third column presents a blended image that combines the generated image with the scribe graph, highlighting the precise control of the model through boundary alignment. This visualization showcases the model’s ability to accurately interpret prompts and maintain robust control over generated images.
@@
In Fig. 11, we present images generated by our model using various prompts. SANA-Sprint showcases comprehensive generation capabilities, including high-fidelity detail rendering, accurate semantic understanding, and reliable text generation, all achieved with only 2-step sampling. In particular, the model efficiently produces high-quality images of 1024 $\times$ 1024 pixels in only 0.24 seconds on an NVIDIA A100 GPU. The samples demonstrate versatility in various scenarios, from in tricate textures and complex compositions to accurate text rendering, highlighting the robust image quality of the model in both artistic and practical tasks.

### v4

@@ 6 Conclusion
In this paper, we introduced SANA-Sprint, an efficient diffusion model for ultra-fast one-step text-to-image generation while preserving multi-step sampling flexibility. By employing a hybrid distillation strategy combining continuous-time consistency distillation (sCM) and latent adversarial distillation (LADD), SANA-Sprint achieves SoTA performance with 7.04 FID and 0.72 GenEval in one step, eliminating step-specific training. This unified step-adaptive model enables high-quality $\times1024$ image generation in only 0.1s on H100, setting a new SoTA in speed-quality tradeoffs.
@@ Appendix A Pseudo Code for Training-Free Transformation from Flow to Trigflow
In this section, we provide a concise implementation of the transformation from a trained flow matching model to a TrigFlow model without requiring additional training. This transformation is based on the theoretical equivalence established in Appendix˜D. The core idea is to first convert the TrigFlow timestep $t_{\texttt{Trig}}$ to its corresponding flow matching timestep $t_{\texttt{FM}}$. Then, the input feature $\boldsymbol{x}_{\texttt{Trig}}$ is scaled accordingly to obtain $\boldsymbol{x}_{\texttt{FM}}$. The output of the flow matching model is then transformed using a linear combination to produce the final TrigFlow output. The following pseudo code implements this transformation efficiently, ensuring consistency between the two formulations.
@@ Appendix B Transformation Algorithm
We present an algorithm for training-free transformation from a flow matching model to its TrigFlow counterpart. Given a noisy sample, its corresponding TrigFlow timestep, and a pre-trained flow matching model, the algorithm computes the equivalent flow matching timestep, rescales the input, and applies a deterministic transformation to obtain the TrigFlow output. The detailed procedure is outlined in Algorithm˜1.
@@ Appendix C Training Algorithm of SANA-Sprint
11:    $\boldsymbol{\psi}\leftarrow\boldsymbol{\psi}-\eta\nabla_{\boldsymbol{\psi}}\mathcal{L}_{\text{adv}}^{D}(\boldsymbol{\psi})$ $\triangleright$ Discriminator step
@@
17:    $r\leftarrow\min(1,\text{Iters}/H)$ $\triangleright$ Tangent warmup
18:    $\boldsymbol{g}\leftarrow-\cos^{2}(t)(\sigma_{d}\boldsymbol{F_{\theta^{-}}}-\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t})-r\cdot\cos(t)\sin(t)(\boldsymbol{x}_{t}+\sigma_{d}\frac{\mathrm{d}\boldsymbol{F_{\theta^{-}}}}{\mathrm{d}t})$ $\triangleright$ JVP rearrangement
19:    $\boldsymbol{g}\leftarrow\boldsymbol{g}/(\|\boldsymbol{g}\|+c)$ $\triangleright$ Tangent normalization
20:    $\mathcal{L}(\boldsymbol{\theta,\phi})\leftarrow\frac{e^{w_{\boldsymbol{\phi}}(t)}}{D}\|\boldsymbol{F_{\theta}}(\frac{\boldsymbol{x}_{t}}{\sigma_{d}},t)-\boldsymbol{F_{\theta^{-}}}(\frac{\boldsymbol{x}_{t}}{\sigma_{d}},t)-\boldsymbol{g}\|_{2}^{2}-w_{\boldsymbol{\phi}}(t)$ $\triangleright$ sCM loss
@@
24:    $\mathcal{L}(\boldsymbol{\theta,\phi})\leftarrow\mathcal{L}(\boldsymbol{\theta,\phi})-\mathbb{E}_{\boldsymbol{x}_{0},s,t}\Big[\sum_{k}D_{\boldsymbol{\psi},k}(\boldsymbol{F}_{\boldsymbol{\theta}_{\texttt{pre}},k}(\hat{\boldsymbol{x}}_{s}^{\boldsymbol{f_{\theta}}},s,\boldsymbol{y}))\Big]$ $\triangleright$ GAN loss
25:    $(\boldsymbol{\theta,\phi})\leftarrow(\boldsymbol{\theta,\phi})-\eta\nabla_{\theta,\phi}\mathcal{L}(\boldsymbol{\theta,\phi})$ $\triangleright$ Generator step
@@
Since $(1-t_{\texttt{FM}})^{2}+t_{\texttt{FM}}^{2}$ is generally not equal to $\cos^{2}(t_{\texttt{Trig}})+\sin^{2}(t_{\texttt{Trig}})=1$ (except when $t_{\texttt{FM}}=0$ or 11), a scale adjustment is needed. To align their scales, we introduce a scale factor function $\lambda(t_{\texttt{FM}})$ that satisfies
@@ Appendix E Full Related Work
Text-to-image generation has experienced transformative advancements in both efficiency and model design. The field gained early traction with Stable Diffusion [59], which set the stage for scalable high-resolution synthesis. A pivotal shift occurred with Diffusion Transformers (DiT)[60], which replaced conventional U-Net architectures with transformer-based designs, unlocking improved scalability and computational efficiency. Building on this innovation, PixArt-$\alpha$[61] demonstrated competitive image quality while slashing training costs to only 10.8% of those required by Stable Diffusion v1.5 [59]. Recent breakthroughs have further pushed the boundaries of compositional generation. Large-scale models like FLUX [21] and Stable Diffusion 3 [38] have scaled up to ultra-high-resolution synthesis and introduced multi-modal capabilities through frameworks such as the Multi-modal Diffusion Transformer ($M$-DiT)[22]. Playground v3[39] achieved state-of-the-art image quality by seamlessly integrating diffusion models with Large Language Models (LLMs)[62], while PixArt-$\Sigma$[37] showcased direct 4K image generation using a compact 0.6B parameter model, emphasizing computational efficiency alongside high-quality outputs. Efficiency-driven innovations have also gained momentum. SANA [32] introduced high-resolution synthesis capabilities through deep compression autoencoding [63] and linear attention mechanisms, enabling deployment on consumer-grade hardware like laptop GPUs. Additionally, advancements in linear attention mechanisms for class-conditional generation [64, 65], diffusion models without attention [66, 67], and cascade structures [68, 69, 70] have further optimized computational requirements while maintaining performance. These developments collectively underscore the field’s rapid evolution toward more accessible, efficient, and versatile text-to-image generation technologies.
@@ F.1 Experimental Setup
We conduct distributed training using PyTorch’s Distributed Data Parallel (DDP) across 32 NVIDIA A100 GPUs on 4 DGX nodes. Our two-phase strategy involves fine-tuning the teacher model with dense time embedding and QK normalization at a learning rate of 2e-5 for 5,000 iterations (global batch size of 1,024), as discussed in Sec.˜3.1. Then, we perform timestep distillation through the proposed framework at a learning rate of 2e-6 with a global batch size of 512 for 20,000 iterations. As Flash Attention JVP kernel support is not available in PyTorch [17], we retain Linear Attention [32] to auto-compute the JVP.
@@ F.2 More Ablations
Fig.˜6 illustrates the process of timestep optimization for inference across 1, 2, and 4 steps, comparing the performance of 0.6B and 1.6B models in terms of FID (top row) and CLIP-Scores (bottom row). The optimization follows a sequential search strategy: first, we determine the optimal $t_{\text{max}}$ for 1-step inference using $\arctan(n/0.5)$, inspired by EDM [28], where $n$ is searched for the maximum timestep. Using this $t_{\text{max}}$, we then search for the intermediate timestep $t_{2nd}$ in 2-step inference. For 4-step inference, the timesteps for the first two steps are fixed to their previously optimized values, while the third ($t_{3rd}$) and fourth timesteps ($t_{4th}$) are searched sequentially. In each case, the x-axis represents the timestep being optimized at the current step, ensuring that earlier steps use their best-found values to maximize overall performance. This hierarchical approach enables efficient timestep selection for multi-step inference settings.
@@
Timestep T $\pi/2,0.0] [$$\arctan(200/0.5),1.3,0.0] [$$\arctan(200/0.5),1.3,1.1,0.6,0.0]$
@@
In the sCM-only experiments, we investigate the impact of different noise distribution parameter settings on model performance. The noise distribution is defined as $t=\arctan{\left(\frac{e^{\tau}}{\sigma_{d}}\right)}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$. Starting from the initial parameters (0.8,1.6)(-0.8,1.6) proposed in sCM [17], we experiment with various mean and standard deviation configurations to evaluate their effects. By tracking FID and CLIP-Score trends over 40k training iterations, we identify $(P_{\text{mean}},P_{\text{std}})=(0.0,1.6)$, represented by the green curve in Fig.˜7, as the optimal setting. This configuration consistently reduces FID while improving CLIP-Score, resulting in superior generation quality and text-image alignment. We also observe that extreme mean values, such as $P_{\text{mean}}=0.6$ or $P_{\text{mean}}=-0.8$, lead to significant training instability and even failure in some cases. Consequently, we adopt $(0.0,1.6)$ as the default parameter setting.
@@
Generative features change with the noise level, offering structured feedback at high noise and texture-related feedback at low noise [5]. We compare the results for different mean and standard deviation settings in $t=\arctan{\left(\frac{e^{\tau}}{\sigma_{d}}\right)}$, where $\tau\sim\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ for LADD discriminator. Building on the optimal mean and standard deviation settings $(0.0,1.6)$ identified for sCM, we further explore the best noise configuration for the LADD’s discriminator. In Fig.˜8 and Tab.˜8, we visualize the probability distributions of $t$ sampled under different mean and standard deviation settings, as well as the corresponding FID and CLIP-Score results when applied in the LADD loss. Based on these analyses, we identify (0.6,1.0)(-0.6,1.0) as the optimal setting, which achieves a more balanced feature distribution across high and low noise levels while maintaining stable training dynamics. Consequently, we adopt (0.6,1.0)(-0.6,1.0) as the default configuration for LADD loss.
@@ F.3 More Qualitative Results
In Fig.˜9, we demonstrate the visualization capabilities of our SANA-Sprint-ControlNet, which efficiently achieves impressive results in only 0.4 seconds using a 2-step generation process, producing high-quality images at a resolution of 1024 $\times$ 1024 pixels. The visualization process begins with an input image, which is processed using a HED detection model to extract the scribe graph. This scribe graph, combined with a given prompt, is used to generate the corresponding image in the second column. The third column presents a blended image that combines the generated image with the scribe graph, highlighting the precise control of the model through boundary alignment. This visualization showcases the model’s ability to accurately interpret prompts and maintain robust control over generated images.
@@
In Fig.˜11, we present images generated by our model using various prompts. SANA-Sprint showcases comprehensive generation capabilities, including high-fidelity detail rendering, accurate semantic understanding, and reliable text generation, all achieved with only 2-step sampling. In particular, the model efficiently produces high-quality images of 1024 $\times$ 1024 pixels in only 0.24 seconds on an NVIDIA A100 GPU. The samples demonstrate versatility in various scenarios, from in tricate textures and complex compositions to accurate text rendering, highlighting the robust image quality of the model in both artistic and practical tasks.